A current of 2.31 A flows through a 1.11 mm diameter copper wire that is connected to a capacitor plate. The current is directed toward the plate. Find the charge transferred to the plate in 9.10 x 10-4 s.
Answer #1

Solution)

Charge transferred to the plate, Q= It

I=2.3 A

Time taken, t=9.10*10^-4 secs

Q= It = 2.3 *(9.10*10^-4)=0.002093 C (Ans)

=====